Which of the following does not represent an integer?
(a) 0 ÷ (– 7) (b) 20 ÷ (– 4) (c) (– 9) ÷ 3 (d) (– 12) ÷ 5

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Option (d) is correct here

0 ÷ (–7) = 0, an integer
20 ÷ (–4) = -5, an integer
(–9) ÷ 3 = -3, an integer
But, (–12) ÷ 5 = -2.4, which is a decimal and not an integer

Thanks